Solve the equation 13z/z+1 = 11-3i zeC [where C denotes the set of complex numbers]
1
Expert's answer
2020-11-08T18:41:55-0500

Let "z=x+iy."

"13x+13iy=(x+iy+1)(11-3i);"

"13x+13iy=11x+11iy+11-3ix+3y-3i;"

"2x-3y+i(3x+2y)=11-3i;"

"\begin{cases}
 2x-3y=11, \\
 3x+2y=-3;
\end{cases}"

"\begin{cases}
 x=\frac{11+3y}{2},\\
 \frac32(11+3y)+2y=-3;
\end{cases}"

"16.5+4.5y+2y=-3;"

"6.5y=-19.5;"

"y=-3;"

"x=1."

"z=1-3i."